OpenStudy (anonymous):

If the set of real roots of the function f(x) is {-2,1,3}, find the set of real roots of f(x^2) and f(x^3-6)

OpenStudy (anonymous):

\[f(x)=(x+2)(x-1)(x-3)\\ f(x^2)=(x^2+2)(x^2-1)(x^2-3)\]

OpenStudy (anonymous):

\[x^2+2=0\] has no real solution \[x^2-1=0\iff x =\pm1\\ x^2-3=0\iff x=\pm\sqrt3\]

OpenStudy (aum):

f(-2) = 0; f(1) = 0; f(3) = 0 (given) Take f(-2) = 0 f(x^3-6) will be zero when x^3 - 6 = -2 or x^3 = 4 or \(\large x = \sqrt[3]{4}\) Take f(1) = 0 f(x^3-6) will be zero when x^3 - 6 = 1 or x^3 = 7 or \(\large x = \sqrt[3]{7}\) Try the case of f(3) = 0 ....